OpenStudy (anonymous):

hey guys i need your help i don't really know how to solve this 3x-y=5 and 2x+y=-10 in substitution way.tnx

OpenStudy (anonymous):

-y = 5 - 3x y = -5 + 3x or y= 3x-5 substituting it to the 2nd equation, we have: 2x + (3x - 5) = -10 5x - 5 = -10 5x = -10 +5 5x = -5 x = -1 y = 3x - 5 y = 3(-1) - 5 y = -3 - 5 y = -8

OpenStudy (anonymous):

thanks you so much

OpenStudy (anonymous):

No prob :))

OpenStudy (anonymous):

substitution means to make one equation equal one of the variables in the second then plug it in so: 3x-y=5 lets solve for y 3x-5=y plug that into the second equation. 2x+y=-10 2x+(3x-5)=-10 5x-5=10 5x=-5 x=-1 then plug that into either of the origonal equations to find y 2x+y=-10 2(-1)+y=-10 -2+y=-10 y=-8 and i realized that someone already finished this but i hope it still helps

OpenStudy (anonymous):

@jdhyde, haha. XD

OpenStudy (anonymous):

ya im not all that fast at typing but i know my math

OpenStudy (anonymous):

yeah really big help thank you for spending your time with me.

OpenStudy (anonymous):

Good for you. :))